OpenStudy (anonymous):

Use the rules of differentiation to find f'(x) for f(x) equal to: 3-4/x^2

OpenStudy (callisto):

\[3-\frac{4}{x^2}\] Is this the equation given?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

\[f'(x)=-(-2)\frac{4}{x^{2+1}}= \frac{8}{x^3}\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (callisto):

welcome :)